# Logical perspectives on learning statistical objects

### v1

@@ 5.2 Bounding the mean width for a derived class in terms of a base class
In this subsection, we will establish connections between combinatorial dimensions of a class of sets or functions and dimensions of its randomization class. Following the approach in (Ben Yaacov, 2009), we will first establish this connection for notions of mean width.
We will show that Rademacher mean width does not increase under averaging. With that in hand, if we are able to bound learnability of $f$ through mean width, the same bound will apply to $\mathbb{E}[f]$. This strategy stems from (Ben Yaacov, 2009, Theorem 4.1), where it was applied to Gaussian mean width.
Assume that $(\Omega,\Sigma,\mu)$ is a probability space, and consider a family $(f_{\omega}:\omega\in\Omega)$ of functions $X\times Y\to[0,1]$ such that for each $x\in X^{n}$ and $v\in{\mathcal{R}}^{n}$, the function $\omega\mapsto h_{f_{\omega}(\bar{x},Y)}(v)$ is measurable. This measurability follows trivially when $Y$ is countable, and will also hold when these functions derive from the randomization class.
We now note a relationship between suprema of expectations and expectations of suprema.
Lemma 3.
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let $(f_{\omega}:\omega\in\Omega)$ be such that $\omega\mapsto h_{f_{\omega}(\bar{x},Y)}(v)$ is measurable. Fix $\bar{x}\in X^{n}$ and $v\in{\mathcal{R}}^{n}$. Then
$h_{\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y)}(v)\leq\mathbb{E}_{\mu}[h_{(f_{\omega}(\bar{x},Y)}(v)].$
Proof.
The supremum of the expectations of a family of functions is at most the expectation of their suprema, so we have
$h_{\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y)}(v)$ $=\sup_{y\in Y}v\cdot(\mathbb{E}_{\mu}[f_{\omega}](\bar{x},y))$
$=\sup_{y\in Y}\mathbb{E}_{\mu}[v\cdot(f_{\omega}(\bar{x},y))]$
$\leq\mathbb{E}_{\mu}\left[\sup_{y\in Y}v\cdot f_{\omega}(\bar{x},y)\right]$
$=\mathbb{E}_{\mu}[h_{f_{\omega}(\bar{x},Y)}(v)].$
which proves the lemma.  [^2] This result is implicit in the proof of (Ben Yaacov, 2009, Theorem 4.1). For a fixed $\bar{x}$, it is stated there that $\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y)\subseteq\mathbb{E}_{\mu}[\overline{\mathrm{Conv}}(f_{\omega}(\bar{x},Y))]$, where the latter expectation is an expectation of convex compact sets. This amounts to saying that for every $\vec{v}\in{\mathcal{R}}^{n}$, $h_{\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y)}(\vec{v})\leq\mathbb{E}_{\mu}[h_{f_{\omega}(\bar{x},Y)}(\vec{v})].$
And we can now make a conclusion about how mean width behaves under expectation.
Lemma 4.
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let $(f_{\omega}:\omega\in\Omega)$ be a family of real-valued functions $f:X\times Y\to[0,1]$ such that for each $(x,y)\in X\times Y$ and $v\in{\mathcal{R}}^{n}$, the functions $\omega\mapsto f_{\omega}(x,y)$ and $\omega\mapsto h_{f_{\omega}(\bar{x},Y)}(v)$ are measurable.
@@
$0pt(\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y),\beta)\leq\mathbb{E}_{\mu}[0pt(f_{\omega}(\bar{x},Y),\beta)].$
Proof.
To prove this, we only need to unfold the definition of $0pt(A,\beta)$ and apply Lemma 3 and Fubini’s Theorem.
$w(\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y),\beta)$ $=\mathbb{E}_{\beta}\left[h_{\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y)}(\vec{b})\right]$
$\leq\mathbb{E}_{\beta}\mathbb{E}_{\mu}[h_{f_{\omega}(\bar{x},Y)}(\vec{b})]$
$=\mathbb{E}_{\mu}\mathbb{E}_{\beta}[h_{f_{\omega}(\bar{x},Y)}(\vec{b})]$
$=\mathbb{E}_{\mu}\left[w(h_{f_{\omega}(\bar{x},Y)},\beta)\right]$
We are now ready to bound the Rademacher mean width of an expectation using the Rademacher mean width of the underlying class, as mentioned in the body of the paper:
Theorem 1 (Pushing a Mean Width Bound through an Expectation).
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let $(f_{\omega}:\omega\in\Omega)$ be a family of real-valued functions $f:X\times Y\to[0,1]$ such that for each $(x,y)\in X\times Y$ and $v\in{\mathcal{R}}^{n}$, the functions $\omega\mapsto f_{\omega}(x,y)$ and $\omega\mapsto h_{f_{\omega}(\bar{x},Y)}(v)$ are measurable.
@@
${\mathcal{R}}_{\mathbb{E}[f]}(n)\leq\sup_{\omega}{\mathcal{R}}_{f_{\omega}}(n).$
Proof.
For each $n$, where $\beta$ is uniformly distributed on $\{-1,1\}^{n}$, it suffices to show that
$\sup_{\bar{x}\in X^{n}}w(\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y),\beta)\leq\sup_{\omega}\sup_{\bar{x}\in X^{n}}w(f_{\omega}(\bar{x},Y),\beta),$
@@
$w(\mathbb{E}_{\mu}[f_{\omega}](\bar{x},Y),\beta)\leq\sup_{\omega}w(f_{\omega}(\bar{x},Y),\beta),$
@@
$\mathbb{E}_{\mu}[w(f_{\omega}(\bar{x},Y),\beta)]\leq\sup_{\omega}w(f_{\omega}(\bar{x},Y),\beta).$
@@ 5.3 Glivenko-Cantelli Bounds through Mean Width
Above we have seen how to estimate how moving to an expectation impacts means width. We will now look at the impact on Glivenko-Cantelli dimension. We can bound the Glivenko-Cantelli dimension with Rademacher mean width. The following is a restatement of (Wainwright, 2019, Theorem 4.10) in terms of $GC$-dimension, using the fact that for any probability measure $\mu$ on $X$ and any $f$, the Rademacher complexity $\frac{1}{n}\mathbb{E}_{\mu^{n}}[w_{\mathcal{R}}(f(\bar{x},Y))]$ is at most $\frac{1}{n}{\mathcal{R}}_{f}(n)$.
Fact 13.
Let $f:X\times Y\to[0,1]$ be a real-valued function, which we view as a hypothesis class on $X$ parametrized by $Y$. For any $\delta>0$ and $n$, then
$GC_{f}\left(\frac{2\cdot{\mathcal{R}}_{f}(n)}{n}+\delta,\exp\left(-\frac{n\delta^{2}}{2}\right)\right)\leq n.$
@@
Lemma 5 (From Rademacher Width of a Base Class to GC of the Expectation class).
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let $(f_{\omega}:\omega\in\Omega)$ be a family of real-valued functions $f:X\times Y\to[0,1]$ such that for each $(x,y)\in X\times Y$ and $v\in{\mathcal{R}}^{n}$, the functions $\omega\mapsto f_{\omega}(x,y)$ and $\omega\mapsto h_{f_{\omega}(\bar{x},Y)}(v)$ are measurable.
For any $\epsilon,\delta>0$, if $N$ is such that for all $n\geq N$, $\frac{{\mathcal{R}}_{f_{\omega}}(n)}{n}\leq\frac{\epsilon}{4}$ for each $\omega\in\Omega$, then
$GC_{\mathbb{E}[f]}(\epsilon,\delta)\leq N+\frac{8}{\epsilon^{2}}\log\frac{1}{\delta}.$
@@
Proof.
Suppose that for all $n\geq N$ and $\omega\in\Omega$, $\frac{{\mathcal{R}}_{f_{\omega}}(n)}{n}\leq\frac{\epsilon}{4}$. Now fix $n\geq N+\frac{8}{\epsilon^{2}}\log\frac{1}{\delta}$, and observe that $\frac{{\mathcal{R}}_{f_{\omega}}(n)}{n}\leq\frac{\epsilon}{4}$ still holds for all $\omega\in\Omega$.
Then by Theorem 1, we see that $n$ is also large enough that $\frac{{\mathcal{R}}_{\mathbb{E}[f]}(n)}{n}\leq\frac{\epsilon}{4}$. Then setting $\gamma=\epsilon-\frac{2{\mathcal{R}}_{\mathbb{E}[f]}(n)}{n}$, our assumption implies that $\gamma\geq\frac{\epsilon}{2}$. Plugging in $\gamma$ for $\delta$ in Fact 13 we have $GC_{\mathbb{E}[f]}\left(\epsilon,\exp\left(-\frac{n\gamma^{2}}{2}\right)\right)\leq n$. As $\gamma\geq\frac{\epsilon}{2}$ and $n\geq\frac{8}{\epsilon^{2}}\log\frac{1}{\delta}$, we have
@@
Thus the conclusion holds. ∎
@@ 5.4 Proof of Theorem 5 in the concept class case
In this setting, we can estimate ${\mathcal{R}}_{f}(n)$ in terms of VC-dimension. For a concept class over $X$, indexed by $Y$, the characteristic function of the class is a function $X\times Y\to\{0,1\}$.
Fact 14 ((Wainwright, 2019, Lemma 4.14 and Equation 4.24)).
Assume that $f$ is the characteristic function of a concept class with VC-dimension at most $d_{f}$. Then for $n\geq 1$,
${\mathcal{R}}_{f}(n)\leq 2\cdot\sqrt{d_{f}\cdot n\cdot\log(n+1)}.$
@@
Recall that ${\mathcal{H}}$ is a $\{0,1\}$-valued class (that is, a concept class) with VC-dimension at most $d$. Putting Fact 14 together with Lemma 5, we see that to bound the GC-dimension of the randomization class of ${\mathcal{H}}$, it suffices to find $N$ large enough that for $n\geq N$, $2\sqrt{\frac{d\log(n+1)}{n}}\leq\frac{\epsilon}{4}$, and add $\frac{1}{\epsilon^{2}}\log\frac{1}{\delta}$. This inequality is equivalent to
@@
$GC_{\mathbb{E}[f]}(\epsilon,\delta)\leq N+\frac{8}{\epsilon^{2}}\log\frac{1}{\delta}=O\left(\frac{1}{\epsilon^{2}}\left(d\log\frac{d}{\epsilon}+\log\frac{1}{\delta}\right)\right).$
By Fact 11, this completes the proof of Theorem 5 in the case of concept classes.
@@ 5.5 Extension to the real-valued case
We now extend to get sample complexity bounds for random objects, but where we start with a class of real-valued functions. Our aim will be:
Theorem 2.
For any $\epsilon,\delta>0$, if $d$ is the $\frac{\epsilon}{50}$ fat-shattering dimension of $f$, then the sample complexity of agnostic PAC learning is bounded by:
$O\left(\frac{d}{\epsilon^{4}}\log^{2}\frac{d}{\epsilon}+\frac{1}{\epsilon^{2}}\log\frac{1}{\delta}\right)$
@@
We will use covering numbers. We first recall unnormalized $\ell_{p}$ norms: the $\ell_{p}$ norm on ${\mathcal{R}}^{n}$ is $\left(\sum_{i=1}^{n}x_{i}^{p}\right)^{1/p}$, while $\ell_{\infty}$ is $\max_{i=1}^{n}x_{i}$.
Definition 12.
For $A\subseteq{\mathcal{R}}^{n}$, $\gamma>0$, and $1\leq p\leq\infty$, we let $\mathcal{N}_{p}(\gamma,A)$ denote the minimum number of $\gamma$-balls in the $\ell_{p}$-metric that cover $A$.
We also let $\mathcal{N}_{p}(\gamma,f,n)$ denote $\sup_{\bar{x}\in X^{n}}\mathcal{N}_{p}(\gamma,f(\bar{x},Y)),$ with $f(\bar{x},Y)$ defined as in Definition 10.
We have defined these for arbitrary $p$, but from now we will use only $p=2$ and $p=\infty$. The relevant relation between them is that for all $x\in{\mathcal{R}}^{n}$, we have $|x|_{2}\leq\sqrt{n}|x|_{\infty}$, so for any $A\subseteq{\mathcal{R}}^{n}$, $\mathcal{N}_{2}(\gamma\sqrt{n},A)\leq\mathcal{N}_{\infty}(\gamma,A)$, and for any $f$ and $n$,
$\mathcal{N}_{2}(\gamma\sqrt{n},f,n)\leq\mathcal{N}_{\infty}(\gamma,f,n).$
@@
Fact 15 (From the proof of (Alon et al., 1997, Lemma 3.5)).
Let $f:X\times Y\to[0,1]$ be a real-valued function, which we view as a hypothesis class on $X$ parametrized by $Y$. Let $d$ be the $\frac{\gamma}{4}$ fat-shattering dimension of $f$. Then
$\mathcal{N}_{\infty}(\gamma,f,n)\leq 2\left(\frac{4n}{\gamma^{2}}\right)^{d\log(2en/d\gamma)}.$
@@
Fact 16 ((Wainwright, 2019, Exercise 5.5)).
@@
Fact 17 ((Wainwright, 2019, Equation 5.36)).
@@
We will concern ourselves with $A\subseteq[0,1]^{n}$, so $D\leq\sqrt{n}$. Thus for $0\leq\gamma\leq 1$, we may plug in $\gamma\sqrt{n}\leq D$, and get
@@
Theorem 3.
Let $f:X\times Y\to[0,1]$, thus $f$ can be considered as a hypothesis class of real valued functions on $X$, indexed by $Y$. For any $\epsilon,\delta>0$, if $d$ is the $\frac{\epsilon}{50}$ fat-shattering dimension of $f$, then
$GC_{\mathbb{E}[f]}(\epsilon,\delta)=O\left(\frac{d}{\epsilon^{4}}\log^{2}\frac{d}{\epsilon}+\frac{1}{\epsilon^{2}}\log\frac{1}{\delta}\right)$
Proof.
By Lemma 5, it suffices to show that there is a constant $C>0$ such that if
@@
$\frac{{\mathcal{R}}_{f}(n)}{n}\leq\frac{\epsilon}{4}.$
@@
$\frac{{\mathcal{R}}_{f}(n)}{n}$ $=\sup_{\bar{x}\in X^{n}}\frac{w_{\mathcal{R}}(f(\bar{x},Y))}{n}$
$\leq\sup_{\bar{x}\in X^{n}}\sqrt{\frac{\pi}{2}}\left(\gamma+2\sqrt{\frac{\log\mathcal{N}_{\infty}(\gamma,f(\bar{x},Y))}{n}}\right)$
$\leq\sqrt{\frac{\pi}{2}}\left(\gamma+2\sqrt{\frac{\log\mathcal{N}_{\infty}(\gamma,f,n)}{n}}\right)$
$=\frac{\epsilon}{8}+\sqrt{\frac{2\pi\log\mathcal{N}_{\infty}(\gamma,f,n)}{n}}.$
@@
$\sqrt{\frac{2\pi\log\mathcal{N}_{\infty}(\gamma,f,n)}{n}}\leq\frac{\epsilon}{8},$
@@
$\frac{\log\mathcal{N}_{\infty}(\gamma,f,n)}{n}\leq\frac{\epsilon^{2}}{128\pi}.$
By Fact 15, we see that for all $n$,
$\log\mathcal{N}_{\infty}(\gamma,f,n)=O\left(d\log\left(\frac{4n}{\gamma^{2}}\right)\log\left(\frac{2en}{d\gamma}\right)\right)=O\left(d\log^{2}\left(\frac{n}{\gamma^{2}}\right)\right)=O\left(d\log^{2}\left(\frac{32\pi n}{\epsilon^{2}}\right)\right).$
@@
$\log\mathcal{N}_{\infty}(\gamma,f,n)\leq Dd\log^{2}\left(\frac{32\pi n}{\epsilon^{2}}\right).$

### v2

@@ 4.2 Bounding the mean width for a derived class in terms of a base class
In this subsection, we will establish connections between combinatorial dimensions of each class of sets or functions in a measurable family and the dimensions of the expectation class of the family. Following the approach in (Ben Yaacov, 2009), we will first establish this connection for notions of mean width.
We will show that Rademacher mean width does not increase under averaging. With that in hand, if we are able to bound learnability of each class in a family ${\mathcal{F}}$ through mean width, the same bound will apply to $E{\mathcal{F}}$. This strategy stems from (Ben Yaacov, 2009, Theorem 4.1), where it was applied to Gaussian mean width.
We can now make a note about how width of the set of vectors induced by a measurable family behaves under expectation.
Lemma 20
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ be a measurable family of hypothesis classes on $X$. Fix $\bar{x}\in X^{n}$ and $\vec{b}\in{\mathcal{R}}^{n}$. Then
$0pt({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\vec{b})\leq\mathbb{E}_{\mu}[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\vec{b})].$
Here recall from Definition 19 that ${\mathcal{H}}_{\omega}(\bar{x},Y)$ is the image of the set of vectors for hypothesis class ${\mathcal{H}}_{\omega}$ at $\bar{x}$, as $y$ ranges over $Y$.
Proof  The supremum of the expectations of a family of functions is at most the expectation of their suprema, so we have
$0pt({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\vec{b})$ $=\sup_{y\in Y}\vec{b}\cdot({\mathbb{E}}{\mathcal{F}}(\bar{x},y))$
$=\sup_{y\in Y}\mathbb{E}_{\mu}[\vec{b}\cdot({\mathcal{H}}_{\omega}(\bar{x},y))]$
$\leq\mathbb{E}_{\mu}\left[\sup_{y\in Y}\vec{b}\cdot{\mathcal{H}}_{\omega}(\bar{x},y)\right]$
$=\mathbb{E}_{\mu}[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\vec{b})].$
which proves the lemma.  [^2] This result is implicit in the proof of (Ben Yaacov, 2009, Theorem 4.1). For a fixed $\bar{x}$, it is stated there that ${\mathbb{E}}{\mathcal{F}}(\bar{x},Y)\subseteq\mathbb{E}_{\mu}[\overline{\mathrm{Conv}}({\mathcal{H}}_{\omega}(\bar{x},Y))]$, where the latter expectation is an expectation of convex compact sets. This amounts to saying that for every $\vec{b}\in{\mathcal{R}}^{n}$, $0pt({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\vec{b})\leq\mathbb{E}_{\mu}[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\vec{b})].$
And we can now extend that statement about width of sets to a statement about mean width.
Lemma 21
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ be a measurable family of hypothesis classes on $X$.
@@
$0pt({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\beta)\leq\mathbb{E}_{\mu}[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\beta)].$
Proof  To prove this, we only need to unfold the definition of $0pt(A,\beta)$ and apply Lemma 20 and Fubini’s Theorem.
$w({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\beta)$ $=\mathbb{E}_{\beta}\left[0pt({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\vec{b})\right]$
$\leq\mathbb{E}_{\beta}\mathbb{E}_{\mu}[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\vec{b})]$
$=\mathbb{E}_{\mu}\mathbb{E}_{\beta}[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\vec{b})]$
$=\mathbb{E}_{\mu}\left[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\beta)\right]$
We are now ready to bound the Rademacher mean width of an expectation using the Rademacher mean width of the underlying class:
Theorem 22 (Pushing a Mean Width Bound through an Expectation)
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ be a measurable family of hypothesis classes on $X$.
@@
${\mathcal{R}}_{{\mathbb{E}}{\mathcal{F}}}(n)\leq\sup_{\omega}{\mathcal{R}}_{{\mathcal{H}}_{\omega}}(n).$
Proof  For each $n$, where $\beta$ is uniformly distributed on $\{-1,1\}^{n}$, it suffices to show that
$\sup_{\bar{x}\in X^{n}}0pt({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\beta)\leq\sup_{\omega}\sup_{\bar{x}\in X^{n}}0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\beta),$
@@
$0pt({\mathbb{E}}{\mathcal{F}}(\bar{x},Y),\beta)\leq\sup_{\omega}0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\beta),$
@@
$\mathbb{E}_{\mu}[0pt({\mathcal{H}}_{\omega}(\bar{x},Y),\beta)]\leq\sup_{\omega}w({\mathcal{H}}_{\omega}(\bar{x},Y),\beta).$
@@ 4.3 Glivenko-Cantelli Bounds through Mean Width
Above we have seen how to estimate how moving to an expectation impacts means width. We will now look at the impact on Glivenko-Cantelli dimension. We can bound the Glivenko-Cantelli dimension with Rademacher mean width. The following is a restatement of (Wainwright, 2019, Theorem 4.10) in terms of $GC$-dimension, using the fact that for any probability measure $\mu$ on $X$ and any hypothesis class ${\mathcal{H}}$ on $X$ parameterized by $Y$, the Rademacher complexity $\frac{1}{n}\mathbb{E}_{\mu^{n}}[w_{\mathcal{R}}({\mathcal{H}}(\bar{x},Y))]$ is at most $\frac{1}{n}{\mathcal{R}}_{\mathcal{H}}(n)$.
Fact 11
Let ${\mathcal{H}}$ be a hypothesis class on $X$ parameterized by $Y$. For any $\delta>0$ and $n$, then
$GC_{\mathcal{H}}\left(\frac{2\cdot{\mathcal{R}}_{\mathcal{H}}(n)}{n}+\delta,\exp\left(-\frac{n\delta^{2}}{2}\right)\right)\leq n.$
@@
Lemma 23 (From Rademacher Width of a Base Class to GC of the Expectation class)
Let $(\Omega,\Sigma,\mu)$ be a probability space, and let ${\mathcal{F}}=({\mathcal{H}}_{\omega}:\omega\in\Omega)$ be a measurable family of hypothesis classes on $X$.
For any $\epsilon,\delta>0$, if $N$ is such that for all $n\geq N$, $\frac{{\mathcal{R}}_{{\mathcal{H}}_{\omega}}(n)}{n}\leq\frac{\epsilon}{4}$ for each $\omega\in\Omega$, then
$GC_{{\mathbb{E}}{\mathcal{F}}}(\epsilon,\delta)\leq N+\frac{8}{\epsilon^{2}}\log\frac{1}{\delta}.$
@@
Proof  Suppose that for all $n\geq N$ and $\omega\in\Omega$, $\frac{{\mathcal{R}}_{{\mathcal{H}}_{\omega}}(n)}{n}\leq\frac{\epsilon}{4}$. Now fix $n\geq N+\frac{8}{\epsilon^{2}}\log\frac{1}{\delta}$, and observe that $\frac{{\mathcal{R}}_{{\mathcal{H}}_{\omega}}(n)}{n}\leq\frac{\epsilon}{4}$ still holds for all $\omega\in\Omega$.
Then by Theorem 22, we see that $n$ is also large enough that $\frac{{\mathcal{R}}_{{\mathbb{E}}{\mathcal{F}}}(n)}{n}\leq\frac{\epsilon}{4}$. Then setting $\gamma=\epsilon-\frac{2{\mathcal{R}}_{{\mathbb{E}}{\mathcal{F}}}(n)}{n}$, our assumption implies that $\gamma\geq\frac{\epsilon}{2}$. Plugging in $\gamma$ for $\delta$ in Fact 11 we have $GC_{{\mathbb{E}}{\mathcal{F}}}\left(\epsilon,\exp\left(-\frac{n\gamma^{2}}{2}\right)\right)\leq n$. As $\gamma\geq\frac{\epsilon}{2}$ and $n\geq\frac{8}{\epsilon^{2}}\log\frac{1}{\delta}$, we have
@@
Thus the conclusion holds.
@@ 4.4 Proof of Theorem 15 in the concept class case
In this setting, we can estimate ${\mathcal{R}}_{\mathcal{C}}(n)$ in terms of VC-dimension, defining ${\mathcal{R}}_{\mathcal{C}}(n)$
Fact 12 (Wainwright (2019, Lemma 4.14 and Equation 4.24))
Assume that ${\mathcal{C}}$ is a concept class with VC-dimension at most $d_{\mathcal{C}}$. Then for $n\geq 1$,
${\mathcal{R}}_{\mathcal{C}}(n)\leq 2\cdot\sqrt{d_{\mathcal{C}}\cdot n\cdot\log(n+1)}.$
@@
Recall that ${\mathcal{F}}=({\mathcal{C}}_{\omega}:\omega\in\Omega)$ is a measurable family where each ${\mathcal{C}}_{\omega}$ is a $\{0,1\}$-valued class (that is, a concept class) with VC-dimension at most $d$. Putting Fact 12 together with Lemma 23, we see that to bound the GC-dimension of the expectation class ${\mathbb{E}}{\mathcal{F}}$, it suffices to find $N$ large enough that for $n\geq N$, $2\sqrt{\frac{d\log(n+1)}{n}}\leq\frac{\epsilon}{4}$, and add $\frac{1}{\epsilon^{2}}\log\frac{1}{\delta}$. This inequality is equivalent to
@@
$GC_{{\mathbb{E}}{\mathcal{F}}}(\epsilon,\delta)\leq N+\frac{8}{\epsilon^{2}}\log\frac{1}{\delta}=O\left(\frac{1}{\epsilon^{2}}\left(d\log\frac{d}{\epsilon}+\log\frac{1}{\delta}\right)\right).$
By Fact 10, this completes the proof of Theorem 15 in the case of concept classes.
@@ 4.5 Extension to the real-valued case
We now extend to get sample complexity bounds for random objects, but where we start with a measurable family of hypothesis classes. Our aim will be:
Theorem 24
For any $\epsilon,\delta>0$, if ${\mathcal{F}}$ is a measurable family of hypothesis classes such that each class ${\mathcal{H}}\in{\mathcal{F}}$ has $\frac{\epsilon}{50}$ fat-shattering dimension at most $d$, the sample complexity of agnostic PAC learning on ${\mathbb{E}}{\mathcal{F}}$ is bounded by:
$O\left(\frac{d}{\epsilon^{4}}\log^{2}\frac{d}{\epsilon}+\frac{1}{\epsilon^{2}}\log\frac{1}{\delta}\right).$
@@
We will use covering numbers. The $\ell_{p}$ norm on ${\mathcal{R}}^{n}$ is $\left(\sum_{i=1}^{n}x_{i}^{p}\right)^{1/p}$, while $\ell_{\infty}$ is $\max_{i=1}^{n}x_{i}$.
Definition 25
For $A\subseteq{\mathcal{R}}^{n}$, $\gamma>0$, and $1\leq p\leq\infty$, we let $\mathcal{N}_{p}(\gamma,A)$, the $\gamma$ covering number of $A$, be the minimum number of $\gamma$-balls in the $\ell_{p}$-metric that cover $A$.
We also let $\mathcal{N}_{p}(\gamma,{\mathcal{H}},n)$ denote $\sup_{\bar{x}\in X^{n}}\mathcal{N}_{p}(\gamma,{\mathcal{H}}(\bar{x},Y)),$ with ${\mathcal{H}}(\bar{x},Y)\subseteq[0,1]^{n}$ defined as in Definition 19.
We have defined these for arbitrary $p$, but from now we will use only $p=2$ and $p=\infty$. The relevant relation between them is that for all $x\in{\mathcal{R}}^{n}$, we have $|x|_{2}\leq\sqrt{n}|x|_{\infty}$, so for any $A\subseteq{\mathcal{R}}^{n}$, $\mathcal{N}_{2}(\gamma\sqrt{n},A)\leq\mathcal{N}_{\infty}(\gamma,A)$, and for any ${\mathcal{H}}$ and $n$,
$\mathcal{N}_{2}(\gamma\sqrt{n},{\mathcal{H}},n)\leq\mathcal{N}_{\infty}(\gamma,{\mathcal{H}},n).$
@@
Fact 13 (From the proof of (Alon et al., 1997, Lemma 3.5))
Let ${\mathcal{H}}$ be a hypothesis class on $X$ parameterized by $Y$. Let $d$ be the $\frac{\gamma}{4}$ fat-shattering dimension of ${\mathcal{H}}$. Then
$\mathcal{N}_{\infty}(\gamma,{\mathcal{H}},n)\leq 2\left(\frac{4n}{\gamma^{2}}\right)^{d\log(2en/d\gamma)}.$
@@
Fact 14 (Wainwright (2019, Exercise 5.5))
@@
Fact 15 (Wainwright (2019, Equation 5.36))
@@
We will concern ourselves with $A\subseteq[0,1]^{n}$, so $D\leq\sqrt{n}$. Thus for $0\leq\gamma\leq 1$, we plug in $\gamma\sqrt{n}\leq D$, and get
@@
Theorem 28
For any $\epsilon,\delta>0$, if ${\mathcal{F}}$ is a measurable family of hypothesis classes such that each class ${\mathcal{H}}\in{\mathcal{F}}$ has $\frac{\epsilon}{50}$ fat-shattering dimension at most $d$, then
$GC_{{\mathbb{E}}{\mathcal{F}}}(\epsilon,\delta)=O\left(\frac{d}{\epsilon^{4}}\log^{2}\frac{d}{\epsilon}+\frac{1}{\epsilon^{2}}\log\frac{1}{\delta}\right)$
Proof  By Lemma 23, it suffices to show that there is a constant $C>0$ such that if
@@
$\frac{{\mathcal{R}}_{\mathcal{H}}(n)}{n}\leq\frac{\epsilon}{4}.$
@@
$\frac{{\mathcal{R}}_{\mathcal{H}}(n)}{n}$ $=\sup_{\bar{x}\in X^{n}}\frac{w_{\mathcal{R}}({\mathcal{H}}(\bar{x},Y))}{n}$
$\leq\sup_{\bar{x}\in X^{n}}\sqrt{\frac{\pi}{2}}\left(\gamma+2\sqrt{\frac{\log\mathcal{N}_{\infty}(\gamma,{\mathcal{H}}(\bar{x},Y))}{n}}\right)$
$\leq\sqrt{\frac{\pi}{2}}\left(\gamma+2\sqrt{\frac{\log\mathcal{N}_{\infty}(\gamma,{\mathcal{H}},n)}{n}}\right)$
$=\frac{\epsilon}{8}+\sqrt{\frac{2\pi\log\mathcal{N}_{\infty}(\gamma,{\mathcal{H}},n)}{n}}.$
@@
$\sqrt{\frac{2\pi\log\mathcal{N}_{\infty}(\gamma,{\mathcal{H}},n)}{n}}\leq\frac{\epsilon}{8},$
@@
$\frac{\log\mathcal{N}_{\infty}(\gamma,{\mathcal{H}},n)}{n}\leq\frac{\epsilon^{2}}{128\pi}.$
By Fact 13, we see that for all $n$,
$\log\mathcal{N}_{\infty}(\gamma,{\mathcal{H}},n)=O\left(d\log\left(\frac{4n}{\gamma^{2}}\right)\log\left(\frac{2en}{d\gamma}\right)\right)=O\left(d\log^{2}\left(\frac{n}{\gamma^{2}}\right)\right)=O\left(d\log^{2}\left(\frac{32\pi n}{\epsilon^{2}}\right)\right).$
@@
$\log\mathcal{N}_{\infty}(\gamma,{\mathcal{H}},n)\leq Dd\log^{2}\left(\frac{32\pi n}{\epsilon^{2}}\right).$